OpenStudy (anonymous):

Write 3 sqrt(9)^5 in exponential notation

OpenStudy (anonymous):

Is it 9^5/3?

OpenStudy (lgbasallote):

deoends...is that \[\large \sqrt[3]{9}\]

OpenStudy (lgbasallote):

i mean \[\LARGE \sqrt[3]{9^5}\]

OpenStudy (radar):

I think you have it right\[9^{5/3}\]

OpenStudy (anonymous):

Its \[3\sqrt(9)^5\]

OpenStudy (lgbasallote):

oh...well then no...3 is outside right? like \[\large 3 \sqrt 9^5\]

OpenStudy (lgbasallote):

or is 3 small like \[\LARGE \sqrt[3]{9^5}\]

OpenStudy (anonymous):

Sorry its, \[(\sqrt[3]{9}) ^{5}\]

OpenStudy (lgbasallote):

well then you're right

OpenStudy (radar):

That is different it would be:\[3(9^{5/2})\]

OpenStudy (lgbasallote):

\[\LARGE \sqrt[a]{b^c} = b^{c/a}\]

OpenStudy (anonymous):

Yay! =) Thank you @radar and @lgbasallote

OpenStudy (radar):

Oh back to the \[9^{5/3}\] lol

OpenStudy (radar):

You are welcome.

OpenStudy (radar):

You had right to begin with and the \[\sqrt[q]{b ^{c}}=b ^{c/a}\] is a good general way to memorize the way it works.